Vocaloading:

Consider the following surface: S is defined by the equation z = x^2 + y^2. Find the unit normal vector at the point (1, 1, 2) on S.

Vocaloading:

@K1ngofPadlet

K1NGofPadlet:

To find the normal vector at a given point on a surface, you can utilize the gradient. In this case, the surface is defined by the equation z = x^2 * y^2. First, find the partial derivatives of z with deference to x and y: ∂z/∂x = 2xy^2 ∂z/∂y = 2yx^2 Now, evaluate these partial derivatives at the point (1, 1, 2): ∂z/∂x(1, 1) = 2(1)(1)^2 = 2 ∂z/∂y(1, 1) = 2(1)(1)^2 = 2 The gradient of z at (1, 1, 2) is then given by the vector (∂z/∂x, ∂z/∂y). Consequently, the vector is N = (2, 2). Now, you just need to normalize this vector to obtain the unit vector. The magnitude of N is √(2^2 + 2^2) = √8 = 2√2. So, the unit vector at (1, 1, 2) is (2/2√2, 2/2√2), which simplifies to (√2/√2, √2/√2). In conclusion, the unit vector at the point (1, 1, 2) on the surface is (√2/√2, √2/√2). I hope that helps.

Vocaloading:

@k1ngofpadlet wrote:
To find the normal vector at a given point on a surface, you can utilize the gradient. In this case, the surface is defined by the equation z = x^2 * y^2. First, find the partial derivatives of z with deference to x and y: ∂z/∂x = 2xy^2 ∂z/∂y = 2yx^2 Now, evaluate these partial derivatives at the point (1, 1, 2): ∂z/∂x(1, 1) = 2(1)(1)^2 = 2 ∂z/∂y(1, 1) = 2(1)(1)^2 = 2 The gradient of z at (1, 1, 2) is then given by the vector (∂z/∂x, ∂z/∂y). Consequently, the vector is N = (2, 2). Now, you just need to normalize this vector to obtain the unit vector. The magnitude of N is √(2^2 + 2^2) = √8 = 2√2. So, the unit vector at (1, 1, 2) is (2/2√2, 2/2√2), which simplifies to (√2/√2, √2/√2). In conclusion, the unit vector at the point (1, 1, 2) on the surface is (√2/√2, √2/√2). I hope that helps.
Can you explain calculus?

K1NGofPadlet:

@vocaloading wrote:
@k1ngofpadlet wrote:
To find the normal vector at a given point on a surface, you can utilize the gradient. In this case, the surface is defined by the equation z = x^2 * y^2. First, find the partial derivatives of z with deference to x and y: ∂z/∂x = 2xy^2 ∂z/∂y = 2yx^2 Now, evaluate these partial derivatives at the point (1, 1, 2): ∂z/∂x(1, 1) = 2(1)(1)^2 = 2 ∂z/∂y(1, 1) = 2(1)(1)^2 = 2 The gradient of z at (1, 1, 2) is then given by the vector (∂z/∂x, ∂z/∂y). Consequently, the vector is N = (2, 2). Now, you just need to normalize this vector to obtain the unit vector. The magnitude of N is √(2^2 + 2^2) = √8 = 2√2. So, the unit vector at (1, 1, 2) is (2/2√2, 2/2√2), which simplifies to (√2/√2, √2/√2). In conclusion, the unit vector at the point (1, 1, 2) on the surface is (√2/√2, √2/√2). I hope that helps.
Can you explain calculus?
Calculus is fundamentally a branch of mathematics that deals with rates of change and the accumulation of quantities. It provides implements for studying how things change and how quantities gradually accumulate. There are two main branches of calculus: differential calculus and integral calculus. Differential calculus fixates on the concept of a derivative, which quantifications the rate at which a quantity changes. It avails us understand the deportment of functions, such as determining whether they are incrementing or decrementing, finding their critical points and local extrema, and analyzing their curvature. Derivatives have sundry applications, from physics to economics, sanctioning us to model and prognosticate the demeanor of systems in a precise way. Integral calculus, on the other hand, deals with the concept of integration. It avails us find the accumulation of quantities, such as finding the area under a curve, calculating the total distance peregrinated by an object, or determining the total magnitude of a substance over time. Integrals are withal fundamental in solving differential equations, which are equations involving derivatives. Together, differential and integral calculus form a puissant toolset for understanding and modeling a wide range of phenomena in fields like physics, engineering, economics, and more. They are often edified as a component of advanced mathematics courses and are essential for pursuing vocations in scientific or technical fields.

Vocaloading:

@k1ngofpadlet wrote:
@vocaloading wrote:
@k1ngofpadlet wrote:
To find the normal vector at a given point on a surface, you can utilize the gradient. In this case, the surface is defined by the equation z = x^2 * y^2. First, find the partial derivatives of z with deference to x and y: ∂z/∂x = 2xy^2 ∂z/∂y = 2yx^2 Now, evaluate these partial derivatives at the point (1, 1, 2): ∂z/∂x(1, 1) = 2(1)(1)^2 = 2 ∂z/∂y(1, 1) = 2(1)(1)^2 = 2 The gradient of z at (1, 1, 2) is then given by the vector (∂z/∂x, ∂z/∂y). Consequently, the vector is N = (2, 2). Now, you just need to normalize this vector to obtain the unit vector. The magnitude of N is √(2^2 + 2^2) = √8 = 2√2. So, the unit vector at (1, 1, 2) is (2/2√2, 2/2√2), which simplifies to (√2/√2, √2/√2). In conclusion, the unit vector at the point (1, 1, 2) on the surface is (√2/√2, √2/√2). I hope that helps.
Can you explain calculus?
Calculus is fundamentally a branch of mathematics that deals with rates of change and the accumulation of quantities. It provides implements for studying how things change and how quantities gradually accumulate. There are two main branches of calculus: differential calculus and integral calculus. Differential calculus fixates on the concept of a derivative, which quantifications the rate at which a quantity changes. It avails us understand the deportment of functions, such as determining whether they are incrementing or decrementing, finding their critical points and local extrema, and analyzing their curvature. Derivatives have sundry applications, from physics to economics, sanctioning us to model and prognosticate the demeanor of systems in a precise way. Integral calculus, on the other hand, deals with the concept of integration. It avails us find the accumulation of quantities, such as finding the area under a curve, calculating the total distance peregrinated by an object, or determining the total magnitude of a substance over time. Integrals are withal fundamental in solving differential equations, which are equations involving derivatives. Together, differential and integral calculus form a puissant toolset for understanding and modeling a wide range of phenomena in fields like physics, engineering, economics, and more. They are often edified as a component of advanced mathematics courses and are essential for pursuing vocations in scientific or technical fields.
Wow your view of calculus is quite detailed

K1NGofPadlet:

@vocaloading wrote:
@k1ngofpadlet wrote:
@vocaloading wrote:
@k1ngofpadlet wrote:
To find the normal vector at a given point on a surface, you can utilize the gradient. In this case, the surface is defined by the equation z = x^2 * y^2. First, find the partial derivatives of z with deference to x and y: ∂z/∂x = 2xy^2 ∂z/∂y = 2yx^2 Now, evaluate these partial derivatives at the point (1, 1, 2): ∂z/∂x(1, 1) = 2(1)(1)^2 = 2 ∂z/∂y(1, 1) = 2(1)(1)^2 = 2 The gradient of z at (1, 1, 2) is then given by the vector (∂z/∂x, ∂z/∂y). Consequently, the vector is N = (2, 2). Now, you just need to normalize this vector to obtain the unit vector. The magnitude of N is √(2^2 + 2^2) = √8 = 2√2. So, the unit vector at (1, 1, 2) is (2/2√2, 2/2√2), which simplifies to (√2/√2, √2/√2). In conclusion, the unit vector at the point (1, 1, 2) on the surface is (√2/√2, √2/√2). I hope that helps.
Can you explain calculus?
Calculus is fundamentally a branch of mathematics that deals with rates of change and the accumulation of quantities. It provides implements for studying how things change and how quantities gradually accumulate. There are two main branches of calculus: differential calculus and integral calculus. Differential calculus fixates on the concept of a derivative, which quantifications the rate at which a quantity changes. It avails us understand the deportment of functions, such as determining whether they are incrementing or decrementing, finding their critical points and local extrema, and analyzing their curvature. Derivatives have sundry applications, from physics to economics, sanctioning us to model and prognosticate the demeanor of systems in a precise way. Integral calculus, on the other hand, deals with the concept of integration. It avails us find the accumulation of quantities, such as finding the area under a curve, calculating the total distance peregrinated by an object, or determining the total magnitude of a substance over time. Integrals are withal fundamental in solving differential equations, which are equations involving derivatives. Together, differential and integral calculus form a puissant toolset for understanding and modeling a wide range of phenomena in fields like physics, engineering, economics, and more. They are often edified as a component of advanced mathematics courses and are essential for pursuing vocations in scientific or technical fields.
Wow your view of calculus is quite detailed
Thank you. I've always found calculus to be a fascinating subject. It's incredible how it enables us to understand and describe the way things change and interact in the world around us. The concepts of derivatives and integrals, for instance, have countless genuine-world applications, from physics and engineering to economics and biology. Plus, the resplendency of it all lies in the elegant mathematical machinery that underpins it. It's like a language of its own, availing us decipher the secrets of the universe.

Vocaloading:

@k1ngofpadlet wrote:
@vocaloading wrote:
@k1ngofpadlet wrote:
@vocaloading wrote:
@k1ngofpadlet wrote:
To find the normal vector at a given point on a surface, you can utilize the gradient. In this case, the surface is defined by the equation z = x^2 * y^2. First, find the partial derivatives of z with deference to x and y: ∂z/∂x = 2xy^2 ∂z/∂y = 2yx^2 Now, evaluate these partial derivatives at the point (1, 1, 2): ∂z/∂x(1, 1) = 2(1)(1)^2 = 2 ∂z/∂y(1, 1) = 2(1)(1)^2 = 2 The gradient of z at (1, 1, 2) is then given by the vector (∂z/∂x, ∂z/∂y). Consequently, the vector is N = (2, 2). Now, you just need to normalize this vector to obtain the unit vector. The magnitude of N is √(2^2 + 2^2) = √8 = 2√2. So, the unit vector at (1, 1, 2) is (2/2√2, 2/2√2), which simplifies to (√2/√2, √2/√2). In conclusion, the unit vector at the point (1, 1, 2) on the surface is (√2/√2, √2/√2). I hope that helps.
Can you explain calculus?
Calculus is fundamentally a branch of mathematics that deals with rates of change and the accumulation of quantities. It provides implements for studying how things change and how quantities gradually accumulate. There are two main branches of calculus: differential calculus and integral calculus. Differential calculus fixates on the concept of a derivative, which quantifications the rate at which a quantity changes. It avails us understand the deportment of functions, such as determining whether they are incrementing or decrementing, finding their critical points and local extrema, and analyzing their curvature. Derivatives have sundry applications, from physics to economics, sanctioning us to model and prognosticate the demeanor of systems in a precise way. Integral calculus, on the other hand, deals with the concept of integration. It avails us find the accumulation of quantities, such as finding the area under a curve, calculating the total distance peregrinated by an object, or determining the total magnitude of a substance over time. Integrals are withal fundamental in solving differential equations, which are equations involving derivatives. Together, differential and integral calculus form a puissant toolset for understanding and modeling a wide range of phenomena in fields like physics, engineering, economics, and more. They are often edified as a component of advanced mathematics courses and are essential for pursuing vocations in scientific or technical fields.
Wow your view of calculus is quite detailed
Thank you. I've always found calculus to be a fascinating subject. It's incredible how it enables us to understand and describe the way things change and interact in the world around us. The concepts of derivatives and integrals, for instance, have countless genuine-world applications, from physics and engineering to economics and biology. Plus, the resplendency of it all lies in the elegant mathematical machinery that underpins it. It's like a language of its own, availing us decipher the secrets of the universe.
Have u ever thought of game developement as a career?

K1NGofPadlet:

@vocaloading wrote:
@k1ngofpadlet wrote:
@vocaloading wrote:
@k1ngofpadlet wrote:
@vocaloading wrote:
@k1ngofpadlet wrote:
To find the normal vector at a given point on a surface, you can utilize the gradient. In this case, the surface is defined by the equation z = x^2 * y^2. First, find the partial derivatives of z with deference to x and y: ∂z/∂x = 2xy^2 ∂z/∂y = 2yx^2 Now, evaluate these partial derivatives at the point (1, 1, 2): ∂z/∂x(1, 1) = 2(1)(1)^2 = 2 ∂z/∂y(1, 1) = 2(1)(1)^2 = 2 The gradient of z at (1, 1, 2) is then given by the vector (∂z/∂x, ∂z/∂y). Consequently, the vector is N = (2, 2). Now, you just need to normalize this vector to obtain the unit vector. The magnitude of N is √(2^2 + 2^2) = √8 = 2√2. So, the unit vector at (1, 1, 2) is (2/2√2, 2/2√2), which simplifies to (√2/√2, √2/√2). In conclusion, the unit vector at the point (1, 1, 2) on the surface is (√2/√2, √2/√2). I hope that helps.
Can you explain calculus?
Calculus is fundamentally a branch of mathematics that deals with rates of change and the accumulation of quantities. It provides implements for studying how things change and how quantities gradually accumulate. There are two main branches of calculus: differential calculus and integral calculus. Differential calculus fixates on the concept of a derivative, which quantifications the rate at which a quantity changes. It avails us understand the deportment of functions, such as determining whether they are incrementing or decrementing, finding their critical points and local extrema, and analyzing their curvature. Derivatives have sundry applications, from physics to economics, sanctioning us to model and prognosticate the demeanor of systems in a precise way. Integral calculus, on the other hand, deals with the concept of integration. It avails us find the accumulation of quantities, such as finding the area under a curve, calculating the total distance peregrinated by an object, or determining the total magnitude of a substance over time. Integrals are withal fundamental in solving differential equations, which are equations involving derivatives. Together, differential and integral calculus form a puissant toolset for understanding and modeling a wide range of phenomena in fields like physics, engineering, economics, and more. They are often edified as a component of advanced mathematics courses and are essential for pursuing vocations in scientific or technical fields.
Wow your view of calculus is quite detailed
Thank you. I've always found calculus to be a fascinating subject. It's incredible how it enables us to understand and describe the way things change and interact in the world around us. The concepts of derivatives and integrals, for instance, have countless genuine-world applications, from physics and engineering to economics and biology. Plus, the resplendency of it all lies in the elegant mathematical machinery that underpins it. It's like a language of its own, availing us decipher the secrets of the universe.
Have u ever thought of game developement as a career?
Absolutely. As an enjoyer of both mathematics and gaming, the conception of being a game designer has certainly crossed my mind. The faculty to Cookieulate my zealousness for calculus with my ingenious instincts to design immersive gaming experiences sounds incredibly enthusing. I can imagine utilizing calculus to engender authentic physics engines, design involute puzzles, and even develop innovative gameplay mechanics. It would be a thrilling way to explore the intersection of art and science while offering people a chance to have fun and be regaled.

K1NGofPadlet:

@k1ngofpadlet wrote:
@vocaloading wrote:
@k1ngofpadlet wrote:
@vocaloading wrote:
@k1ngofpadlet wrote:
@vocaloading wrote:
@k1ngofpadlet wrote:
To find the normal vector at a given point on a surface, you can utilize the gradient. In this case, the surface is defined by the equation z = x^2 * y^2. First, find the partial derivatives of z with deference to x and y: ∂z/∂x = 2xy^2 ∂z/∂y = 2yx^2 Now, evaluate these partial derivatives at the point (1, 1, 2): ∂z/∂x(1, 1) = 2(1)(1)^2 = 2 ∂z/∂y(1, 1) = 2(1)(1)^2 = 2 The gradient of z at (1, 1, 2) is then given by the vector (∂z/∂x, ∂z/∂y). Consequently, the vector is N = (2, 2). Now, you just need to normalize this vector to obtain the unit vector. The magnitude of N is √(2^2 + 2^2) = √8 = 2√2. So, the unit vector at (1, 1, 2) is (2/2√2, 2/2√2), which simplifies to (√2/√2, √2/√2). In conclusion, the unit vector at the point (1, 1, 2) on the surface is (√2/√2, √2/√2). I hope that helps.
Can you explain calculus?
Calculus is fundamentally a branch of mathematics that deals with rates of change and the accumulation of quantities. It provides implements for studying how things change and how quantities gradually accumulate. There are two main branches of calculus: differential calculus and integral calculus. Differential calculus fixates on the concept of a derivative, which quantifications the rate at which a quantity changes. It avails us understand the deportment of functions, such as determining whether they are incrementing or decrementing, finding their critical points and local extrema, and analyzing their curvature. Derivatives have sundry applications, from physics to economics, sanctioning us to model and prognosticate the demeanor of systems in a precise way. Integral calculus, on the other hand, deals with the concept of integration. It avails us find the accumulation of quantities, such as finding the area under a curve, calculating the total distance peregrinated by an object, or determining the total magnitude of a substance over time. Integrals are withal fundamental in solving differential equations, which are equations involving derivatives. Together, differential and integral calculus form a puissant toolset for understanding and modeling a wide range of phenomena in fields like physics, engineering, economics, and more. They are often edified as a component of advanced mathematics courses and are essential for pursuing vocations in scientific or technical fields.
Wow your view of calculus is quite detailed
Thank you. I've always found calculus to be a fascinating subject. It's incredible how it enables us to understand and describe the way things change and interact in the world around us. The concepts of derivatives and integrals, for instance, have countless genuine-world applications, from physics and engineering to economics and biology. Plus, the resplendency of it all lies in the elegant mathematical machinery that underpins it. It's like a language of its own, availing us decipher the secrets of the universe.
Have u ever thought of game developement as a career?
Absolutely. As an enjoyer of both mathematics and gaming, the conception of being a game designer has certainly crossed my mind. The faculty to Cookieulate my zealousness for calculus with my ingenious instincts to design immersive gaming experiences sounds incredibly enthusing. I can imagine utilizing calculus to engender authentic physics engines, design involute puzzles, and even develop innovative gameplay mechanics. It would be a thrilling way to explore the intersection of art and science while offering people a chance to have fun and be regaled.
Unfortunately, I'm not the very BEST. So I didn't get into any game designing businesses and on top of that, i didn't want to move away from my parents and my girlfriend in case they needed me, you know?

Vocaloading:

@k1ngofpadlet wrote:
@k1ngofpadlet wrote:
@vocaloading wrote:
@k1ngofpadlet wrote:
@vocaloading wrote:
@k1ngofpadlet wrote:
@vocaloading wrote:
@k1ngofpadlet wrote:
To find the normal vector at a given point on a surface, you can utilize the gradient. In this case, the surface is defined by the equation z = x^2 * y^2. First, find the partial derivatives of z with deference to x and y: ∂z/∂x = 2xy^2 ∂z/∂y = 2yx^2 Now, evaluate these partial derivatives at the point (1, 1, 2): ∂z/∂x(1, 1) = 2(1)(1)^2 = 2 ∂z/∂y(1, 1) = 2(1)(1)^2 = 2 The gradient of z at (1, 1, 2) is then given by the vector (∂z/∂x, ∂z/∂y). Consequently, the vector is N = (2, 2). Now, you just need to normalize this vector to obtain the unit vector. The magnitude of N is √(2^2 + 2^2) = √8 = 2√2. So, the unit vector at (1, 1, 2) is (2/2√2, 2/2√2), which simplifies to (√2/√2, √2/√2). In conclusion, the unit vector at the point (1, 1, 2) on the surface is (√2/√2, √2/√2). I hope that helps.
Can you explain calculus?
Calculus is fundamentally a branch of mathematics that deals with rates of change and the accumulation of quantities. It provides implements for studying how things change and how quantities gradually accumulate. There are two main branches of calculus: differential calculus and integral calculus. Differential calculus fixates on the concept of a derivative, which quantifications the rate at which a quantity changes. It avails us understand the deportment of functions, such as determining whether they are incrementing or decrementing, finding their critical points and local extrema, and analyzing their curvature. Derivatives have sundry applications, from physics to economics, sanctioning us to model and prognosticate the demeanor of systems in a precise way. Integral calculus, on the other hand, deals with the concept of integration. It avails us find the accumulation of quantities, such as finding the area under a curve, calculating the total distance peregrinated by an object, or determining the total magnitude of a substance over time. Integrals are withal fundamental in solving differential equations, which are equations involving derivatives. Together, differential and integral calculus form a puissant toolset for understanding and modeling a wide range of phenomena in fields like physics, engineering, economics, and more. They are often edified as a component of advanced mathematics courses and are essential for pursuing vocations in scientific or technical fields.
Wow your view of calculus is quite detailed
Thank you. I've always found calculus to be a fascinating subject. It's incredible how it enables us to understand and describe the way things change and interact in the world around us. The concepts of derivatives and integrals, for instance, have countless genuine-world applications, from physics and engineering to economics and biology. Plus, the resplendency of it all lies in the elegant mathematical machinery that underpins it. It's like a language of its own, availing us decipher the secrets of the universe.
Have u ever thought of game developement as a career?
Absolutely. As an enjoyer of both mathematics and gaming, the conception of being a game designer has certainly crossed my mind. The faculty to Cookieulate my zealousness for calculus with my ingenious instincts to design immersive gaming experiences sounds incredibly enthusing. I can imagine utilizing calculus to engender authentic physics engines, design involute puzzles, and even develop innovative gameplay mechanics. It would be a thrilling way to explore the intersection of art and science while offering people a chance to have fun and be regaled.
Unfortunately, I'm not the very BEST. So I didn't get into any game designing businesses and on top of that, i didn't want to move away from my parents and my girlfriend in case they needed me, you know?
Yeah felt

Vocaloading:

@vocaloading wrote:
@k1ngofpadlet wrote:
@k1ngofpadlet wrote:
@vocaloading wrote:
@k1ngofpadlet wrote:
@vocaloading wrote:
@k1ngofpadlet wrote:
@vocaloading wrote:
@k1ngofpadlet wrote:
To find the normal vector at a given point on a surface, you can utilize the gradient. In this case, the surface is defined by the equation z = x^2 * y^2. First, find the partial derivatives of z with deference to x and y: ∂z/∂x = 2xy^2 ∂z/∂y = 2yx^2 Now, evaluate these partial derivatives at the point (1, 1, 2): ∂z/∂x(1, 1) = 2(1)(1)^2 = 2 ∂z/∂y(1, 1) = 2(1)(1)^2 = 2 The gradient of z at (1, 1, 2) is then given by the vector (∂z/∂x, ∂z/∂y). Consequently, the vector is N = (2, 2). Now, you just need to normalize this vector to obtain the unit vector. The magnitude of N is √(2^2 + 2^2) = √8 = 2√2. So, the unit vector at (1, 1, 2) is (2/2√2, 2/2√2), which simplifies to (√2/√2, √2/√2). In conclusion, the unit vector at the point (1, 1, 2) on the surface is (√2/√2, √2/√2). I hope that helps.
Can you explain calculus?
Calculus is fundamentally a branch of mathematics that deals with rates of change and the accumulation of quantities. It provides implements for studying how things change and how quantities gradually accumulate. There are two main branches of calculus: differential calculus and integral calculus. Differential calculus fixates on the concept of a derivative, which quantifications the rate at which a quantity changes. It avails us understand the deportment of functions, such as determining whether they are incrementing or decrementing, finding their critical points and local extrema, and analyzing their curvature. Derivatives have sundry applications, from physics to economics, sanctioning us to model and prognosticate the demeanor of systems in a precise way. Integral calculus, on the other hand, deals with the concept of integration. It avails us find the accumulation of quantities, such as finding the area under a curve, calculating the total distance peregrinated by an object, or determining the total magnitude of a substance over time. Integrals are withal fundamental in solving differential equations, which are equations involving derivatives. Together, differential and integral calculus form a puissant toolset for understanding and modeling a wide range of phenomena in fields like physics, engineering, economics, and more. They are often edified as a component of advanced mathematics courses and are essential for pursuing vocations in scientific or technical fields.
Wow your view of calculus is quite detailed
Thank you. I've always found calculus to be a fascinating subject. It's incredible how it enables us to understand and describe the way things change and interact in the world around us. The concepts of derivatives and integrals, for instance, have countless genuine-world applications, from physics and engineering to economics and biology. Plus, the resplendency of it all lies in the elegant mathematical machinery that underpins it. It's like a language of its own, availing us decipher the secrets of the universe.
Have u ever thought of game developement as a career?
Absolutely. As an enjoyer of both mathematics and gaming, the conception of being a game designer has certainly crossed my mind. The faculty to Cookieulate my zealousness for calculus with my ingenious instincts to design immersive gaming experiences sounds incredibly enthusing. I can imagine utilizing calculus to engender authentic physics engines, design involute puzzles, and even develop innovative gameplay mechanics. It would be a thrilling way to explore the intersection of art and science while offering people a chance to have fun and be regaled.
Unfortunately, I'm not the very BEST. So I didn't get into any game designing businesses and on top of that, i didn't want to move away from my parents and my girlfriend in case they needed me, you know?
Yeah felt
Are you with Élena or Lisa? I can't remember tbh lol

K1NGofPadlet:

@vocaloading wrote:
@vocaloading wrote:
@k1ngofpadlet wrote:
@k1ngofpadlet wrote:
@vocaloading wrote:
@k1ngofpadlet wrote:
@vocaloading wrote:
@k1ngofpadlet wrote:
@vocaloading wrote:
@k1ngofpadlet wrote:
To find the normal vector at a given point on a surface, you can utilize the gradient. In this case, the surface is defined by the equation z = x^2 * y^2. First, find the partial derivatives of z with deference to x and y: ∂z/∂x = 2xy^2 ∂z/∂y = 2yx^2 Now, evaluate these partial derivatives at the point (1, 1, 2): ∂z/∂x(1, 1) = 2(1)(1)^2 = 2 ∂z/∂y(1, 1) = 2(1)(1)^2 = 2 The gradient of z at (1, 1, 2) is then given by the vector (∂z/∂x, ∂z/∂y). Consequently, the vector is N = (2, 2). Now, you just need to normalize this vector to obtain the unit vector. The magnitude of N is √(2^2 + 2^2) = √8 = 2√2. So, the unit vector at (1, 1, 2) is (2/2√2, 2/2√2), which simplifies to (√2/√2, √2/√2). In conclusion, the unit vector at the point (1, 1, 2) on the surface is (√2/√2, √2/√2). I hope that helps.
Can you explain calculus?
Calculus is fundamentally a branch of mathematics that deals with rates of change and the accumulation of quantities. It provides implements for studying how things change and how quantities gradually accumulate. There are two main branches of calculus: differential calculus and integral calculus. Differential calculus fixates on the concept of a derivative, which quantifications the rate at which a quantity changes. It avails us understand the deportment of functions, such as determining whether they are incrementing or decrementing, finding their critical points and local extrema, and analyzing their curvature. Derivatives have sundry applications, from physics to economics, sanctioning us to model and prognosticate the demeanor of systems in a precise way. Integral calculus, on the other hand, deals with the concept of integration. It avails us find the accumulation of quantities, such as finding the area under a curve, calculating the total distance peregrinated by an object, or determining the total magnitude of a substance over time. Integrals are withal fundamental in solving differential equations, which are equations involving derivatives. Together, differential and integral calculus form a puissant toolset for understanding and modeling a wide range of phenomena in fields like physics, engineering, economics, and more. They are often edified as a component of advanced mathematics courses and are essential for pursuing vocations in scientific or technical fields.
Wow your view of calculus is quite detailed
Thank you. I've always found calculus to be a fascinating subject. It's incredible how it enables us to understand and describe the way things change and interact in the world around us. The concepts of derivatives and integrals, for instance, have countless genuine-world applications, from physics and engineering to economics and biology. Plus, the resplendency of it all lies in the elegant mathematical machinery that underpins it. It's like a language of its own, availing us decipher the secrets of the universe.
Have u ever thought of game developement as a career?
Absolutely. As an enjoyer of both mathematics and gaming, the conception of being a game designer has certainly crossed my mind. The faculty to Cookieulate my zealousness for calculus with my ingenious instincts to design immersive gaming experiences sounds incredibly enthusing. I can imagine utilizing calculus to engender authentic physics engines, design involute puzzles, and even develop innovative gameplay mechanics. It would be a thrilling way to explore the intersection of art and science while offering people a chance to have fun and be regaled.
Unfortunately, I'm not the very BEST. So I didn't get into any game designing businesses and on top of that, i didn't want to move away from my parents and my girlfriend in case they needed me, you know?
Yeah felt
Are you with Élena or Lisa? I can't remember tbh lol
Currently, Élena. Are you from Padlet?

Vocaloading:

@k1ngofpadlet wrote:
@vocaloading wrote:
@vocaloading wrote:
@k1ngofpadlet wrote:
@k1ngofpadlet wrote:
@vocaloading wrote:
@k1ngofpadlet wrote:
@vocaloading wrote:
@k1ngofpadlet wrote:
@vocaloading wrote:
@k1ngofpadlet wrote:
To find the normal vector at a given point on a surface, you can utilize the gradient. In this case, the surface is defined by the equation z = x^2 * y^2. First, find the partial derivatives of z with deference to x and y: ∂z/∂x = 2xy^2 ∂z/∂y = 2yx^2 Now, evaluate these partial derivatives at the point (1, 1, 2): ∂z/∂x(1, 1) = 2(1)(1)^2 = 2 ∂z/∂y(1, 1) = 2(1)(1)^2 = 2 The gradient of z at (1, 1, 2) is then given by the vector (∂z/∂x, ∂z/∂y). Consequently, the vector is N = (2, 2). Now, you just need to normalize this vector to obtain the unit vector. The magnitude of N is √(2^2 + 2^2) = √8 = 2√2. So, the unit vector at (1, 1, 2) is (2/2√2, 2/2√2), which simplifies to (√2/√2, √2/√2). In conclusion, the unit vector at the point (1, 1, 2) on the surface is (√2/√2, √2/√2). I hope that helps.
Can you explain calculus?
Calculus is fundamentally a branch of mathematics that deals with rates of change and the accumulation of quantities. It provides implements for studying how things change and how quantities gradually accumulate. There are two main branches of calculus: differential calculus and integral calculus. Differential calculus fixates on the concept of a derivative, which quantifications the rate at which a quantity changes. It avails us understand the deportment of functions, such as determining whether they are incrementing or decrementing, finding their critical points and local extrema, and analyzing their curvature. Derivatives have sundry applications, from physics to economics, sanctioning us to model and prognosticate the demeanor of systems in a precise way. Integral calculus, on the other hand, deals with the concept of integration. It avails us find the accumulation of quantities, such as finding the area under a curve, calculating the total distance peregrinated by an object, or determining the total magnitude of a substance over time. Integrals are withal fundamental in solving differential equations, which are equations involving derivatives. Together, differential and integral calculus form a puissant toolset for understanding and modeling a wide range of phenomena in fields like physics, engineering, economics, and more. They are often edified as a component of advanced mathematics courses and are essential for pursuing vocations in scientific or technical fields.
Wow your view of calculus is quite detailed
Thank you. I've always found calculus to be a fascinating subject. It's incredible how it enables us to understand and describe the way things change and interact in the world around us. The concepts of derivatives and integrals, for instance, have countless genuine-world applications, from physics and engineering to economics and biology. Plus, the resplendency of it all lies in the elegant mathematical machinery that underpins it. It's like a language of its own, availing us decipher the secrets of the universe.
Have u ever thought of game developement as a career?
Absolutely. As an enjoyer of both mathematics and gaming, the conception of being a game designer has certainly crossed my mind. The faculty to Cookieulate my zealousness for calculus with my ingenious instincts to design immersive gaming experiences sounds incredibly enthusing. I can imagine utilizing calculus to engender authentic physics engines, design involute puzzles, and even develop innovative gameplay mechanics. It would be a thrilling way to explore the intersection of art and science while offering people a chance to have fun and be regaled.
Unfortunately, I'm not the very BEST. So I didn't get into any game designing businesses and on top of that, i didn't want to move away from my parents and my girlfriend in case they needed me, you know?
Yeah felt
Are you with Élena or Lisa? I can't remember tbh lol
Currently, Élena. Are you from Padlet?
Yeah I'm Travis from Emily's chat. Veronica's ex?

K1NGofPadlet:

@vocaloading wrote:
@k1ngofpadlet wrote:
@vocaloading wrote:
@vocaloading wrote:
@k1ngofpadlet wrote:
@k1ngofpadlet wrote:
@vocaloading wrote:
@k1ngofpadlet wrote:
@vocaloading wrote:
@k1ngofpadlet wrote:
@vocaloading wrote:
@k1ngofpadlet wrote:
To find the normal vector at a given point on a surface, you can utilize the gradient. In this case, the surface is defined by the equation z = x^2 * y^2. First, find the partial derivatives of z with deference to x and y: ∂z/∂x = 2xy^2 ∂z/∂y = 2yx^2 Now, evaluate these partial derivatives at the point (1, 1, 2): ∂z/∂x(1, 1) = 2(1)(1)^2 = 2 ∂z/∂y(1, 1) = 2(1)(1)^2 = 2 The gradient of z at (1, 1, 2) is then given by the vector (∂z/∂x, ∂z/∂y). Consequently, the vector is N = (2, 2). Now, you just need to normalize this vector to obtain the unit vector. The magnitude of N is √(2^2 + 2^2) = √8 = 2√2. So, the unit vector at (1, 1, 2) is (2/2√2, 2/2√2), which simplifies to (√2/√2, √2/√2). In conclusion, the unit vector at the point (1, 1, 2) on the surface is (√2/√2, √2/√2). I hope that helps.
Can you explain calculus?
Calculus is fundamentally a branch of mathematics that deals with rates of change and the accumulation of quantities. It provides implements for studying how things change and how quantities gradually accumulate. There are two main branches of calculus: differential calculus and integral calculus. Differential calculus fixates on the concept of a derivative, which quantifications the rate at which a quantity changes. It avails us understand the deportment of functions, such as determining whether they are incrementing or decrementing, finding their critical points and local extrema, and analyzing their curvature. Derivatives have sundry applications, from physics to economics, sanctioning us to model and prognosticate the demeanor of systems in a precise way. Integral calculus, on the other hand, deals with the concept of integration. It avails us find the accumulation of quantities, such as finding the area under a curve, calculating the total distance peregrinated by an object, or determining the total magnitude of a substance over time. Integrals are withal fundamental in solving differential equations, which are equations involving derivatives. Together, differential and integral calculus form a puissant toolset for understanding and modeling a wide range of phenomena in fields like physics, engineering, economics, and more. They are often edified as a component of advanced mathematics courses and are essential for pursuing vocations in scientific or technical fields.
Wow your view of calculus is quite detailed
Thank you. I've always found calculus to be a fascinating subject. It's incredible how it enables us to understand and describe the way things change and interact in the world around us. The concepts of derivatives and integrals, for instance, have countless genuine-world applications, from physics and engineering to economics and biology. Plus, the resplendency of it all lies in the elegant mathematical machinery that underpins it. It's like a language of its own, availing us decipher the secrets of the universe.
Have u ever thought of game developement as a career?
Absolutely. As an enjoyer of both mathematics and gaming, the conception of being a game designer has certainly crossed my mind. The faculty to Cookieulate my zealousness for calculus with my ingenious instincts to design immersive gaming experiences sounds incredibly enthusing. I can imagine utilizing calculus to engender authentic physics engines, design involute puzzles, and even develop innovative gameplay mechanics. It would be a thrilling way to explore the intersection of art and science while offering people a chance to have fun and be regaled.
Unfortunately, I'm not the very BEST. So I didn't get into any game designing businesses and on top of that, i didn't want to move away from my parents and my girlfriend in case they needed me, you know?
Yeah felt
Are you with Élena or Lisa? I can't remember tbh lol
Currently, Élena. Are you from Padlet?
Yeah I'm Travis from Emily's chat. Veronica's ex?
Oh I remember now.

Vocaloading:

Yep

BACKSTREETBOYS3:

This went to helping someone to an ex riveraly?

Vocaloading:

@backstreetboys3 wrote:
This went to helping someone to an ex riveraly?
Huh?

K1NGofPadlet:

@backstreetboys3 wrote:
This went to helping someone to an ex riveraly?
I don't remember asking you to chime in.

Vocaloading:

@k1ngofpadlet wrote:
@backstreetboys3 wrote:
This went to helping someone to an ex riveraly?
I don't remember asking you to chime in.
Chill dude hes fine ig lol

BACKSTREETBOYS3:

Oh my apologies my notifications were blowing up

BACKSTREETBOYS3:

I'm a girl

BACKSTREETBOYS3:

😁

Vocaloading:

@backstreetboys3 wrote:
I'm a girl
OH MY GAWD

Vocaloading:

@backstreetboys3 wrote:
Oh my apologies my notifications were blowing up
SO SO sorry dont hate me

BACKSTREETBOYS3:

Haha it's fine wanna talk in private?

K1NGofPadlet:

@backstreetboys3 wrote:
Haha it's fine wanna talk in private?
About what?

BACKSTREETBOYS3:

I just want to talk to him/her

Vocaloading:

@backstreetboys3 wrote:
I just want to talk to him/her
I am he/him

Vocaloading:

@K1ngofpadlet I'm sorry man another one up: Solve the following equation for x: 3x^2 + 5x - 2 = 0

BACKSTREETBOYS3:

Y'all stop blowing my notifications up

BACKSTREETBOYS3:

Or my phone will blow up

Vocaloading:

@backstreetboys3 wrote:
Or my phone will blow up
Uhhh

BACKSTREETBOYS3:

Sarcasm

K1NGofPadlet:

@backstreetboys3 wrote:
Or my phone will blow up
You just delete your comments and it'll go away..

K1NGofPadlet:

@vocaloading wrote:
@K1ngofpadlet I'm sorry man another one up: Solve the following equation for x: 3x^2 + 5x - 2 = 0
You're good. Utilize the quadratic formula. The quadratic formula states that for any equation in the form ax^2 + bx + c = 0, the solutions can be found utilizing the formula: x = (-b ± √(b^2 - 4ac)) / (2a) In this case, you have a = 3, b = 5, and c = -2. Plugging these values into the formula, you get: x = (-5 ± √(5^2 - 4*3*(-2))) / (2*3) After simplifying, you have: x = (-5 ± √(25 + 24)) / 6 x = (-5 ± √49) / 6 x = (-5 ± 7) / 6 Now you have two possibilities: 1. x = (-5 + 7) / 6 = 2 / 6 = 1/3 2. x = (-5 - 7) / 6 = -12 / 6 = -2 So the solutions to the equation are x = 1/3 and x = -2.

Vocaloading:

@k1ngofpadlet wrote:
@vocaloading wrote:
@K1ngofpadlet I'm sorry man another one up: Solve the following equation for x: 3x^2 + 5x - 2 = 0
You're good. Utilize the quadratic formula. The quadratic formula states that for any equation in the form ax^2 + bx + c = 0, the solutions can be found utilizing the formula: x = (-b ± √(b^2 - 4ac)) / (2a) In this case, you have a = 3, b = 5, and c = -2. Plugging these values into the formula, you get: x = (-5 ± √(5^2 - 4*3*(-2))) / (2*3) After simplifying, you have: x = (-5 ± √(25 + 24)) / 6 x = (-5 ± √49) / 6 x = (-5 ± 7) / 6 Now you have two possibilities: 1. x = (-5 + 7) / 6 = 2 / 6 = 1/3 2. x = (-5 - 7) / 6 = -12 / 6 = -2 So the solutions to the equation are x = 1/3 and x = -2.
Bro idk how u do it

K1NGofPadlet:

@vocaloading wrote:
@k1ngofpadlet wrote:
@vocaloading wrote:
@K1ngofpadlet I'm sorry man another one up: Solve the following equation for x: 3x^2 + 5x - 2 = 0
You're good. Utilize the quadratic formula. The quadratic formula states that for any equation in the form ax^2 + bx + c = 0, the solutions can be found utilizing the formula: x = (-b ± √(b^2 - 4ac)) / (2a) In this case, you have a = 3, b = 5, and c = -2. Plugging these values into the formula, you get: x = (-5 ± √(5^2 - 4*3*(-2))) / (2*3) After simplifying, you have: x = (-5 ± √(25 + 24)) / 6 x = (-5 ± √49) / 6 x = (-5 ± 7) / 6 Now you have two possibilities: 1. x = (-5 + 7) / 6 = 2 / 6 = 1/3 2. x = (-5 - 7) / 6 = -12 / 6 = -2 So the solutions to the equation are x = 1/3 and x = -2.
Bro idk how u do it
Good teacher.

Vocaloading:

@k1ngofpadlet wrote:
@vocaloading wrote:
@k1ngofpadlet wrote:
@vocaloading wrote:
@K1ngofpadlet I'm sorry man another one up: Solve the following equation for x: 3x^2 + 5x - 2 = 0
You're good. Utilize the quadratic formula. The quadratic formula states that for any equation in the form ax^2 + bx + c = 0, the solutions can be found utilizing the formula: x = (-b ± √(b^2 - 4ac)) / (2a) In this case, you have a = 3, b = 5, and c = -2. Plugging these values into the formula, you get: x = (-5 ± √(5^2 - 4*3*(-2))) / (2*3) After simplifying, you have: x = (-5 ± √(25 + 24)) / 6 x = (-5 ± √49) / 6 x = (-5 ± 7) / 6 Now you have two possibilities: 1. x = (-5 + 7) / 6 = 2 / 6 = 1/3 2. x = (-5 - 7) / 6 = -12 / 6 = -2 So the solutions to the equation are x = 1/3 and x = -2.
Bro idk how u do it
Good teacher.
So how is V?

K1NGofPadlet:

@vocaloading wrote:
@k1ngofpadlet wrote:
@vocaloading wrote:
@k1ngofpadlet wrote:
@vocaloading wrote:
@K1ngofpadlet I'm sorry man another one up: Solve the following equation for x: 3x^2 + 5x - 2 = 0
You're good. Utilize the quadratic formula. The quadratic formula states that for any equation in the form ax^2 + bx + c = 0, the solutions can be found utilizing the formula: x = (-b ± √(b^2 - 4ac)) / (2a) In this case, you have a = 3, b = 5, and c = -2. Plugging these values into the formula, you get: x = (-5 ± √(5^2 - 4*3*(-2))) / (2*3) After simplifying, you have: x = (-5 ± √(25 + 24)) / 6 x = (-5 ± √49) / 6 x = (-5 ± 7) / 6 Now you have two possibilities: 1. x = (-5 + 7) / 6 = 2 / 6 = 1/3 2. x = (-5 - 7) / 6 = -12 / 6 = -2 So the solutions to the equation are x = 1/3 and x = -2.
Bro idk how u do it
Good teacher.
So how is V?
She's pretty good. She typically hangs out on Johnson's Padlet.

Vocaloading:

@k1ngofpadlet wrote:
@vocaloading wrote:
@k1ngofpadlet wrote:
@vocaloading wrote:
@k1ngofpadlet wrote:
@vocaloading wrote:
@K1ngofpadlet I'm sorry man another one up: Solve the following equation for x: 3x^2 + 5x - 2 = 0
You're good. Utilize the quadratic formula. The quadratic formula states that for any equation in the form ax^2 + bx + c = 0, the solutions can be found utilizing the formula: x = (-b ± √(b^2 - 4ac)) / (2a) In this case, you have a = 3, b = 5, and c = -2. Plugging these values into the formula, you get: x = (-5 ± √(5^2 - 4*3*(-2))) / (2*3) After simplifying, you have: x = (-5 ± √(25 + 24)) / 6 x = (-5 ± √49) / 6 x = (-5 ± 7) / 6 Now you have two possibilities: 1. x = (-5 + 7) / 6 = 2 / 6 = 1/3 2. x = (-5 - 7) / 6 = -12 / 6 = -2 So the solutions to the equation are x = 1/3 and x = -2.
Bro idk how u do it
Good teacher.
So how is V?
She's pretty good. She typically hangs out on Johnson's Padlet.
makes sense

K1NGofPadlet:

@vocaloading wrote:
@k1ngofpadlet wrote:
@vocaloading wrote:
@k1ngofpadlet wrote:
@vocaloading wrote:
@k1ngofpadlet wrote:
@vocaloading wrote:
@K1ngofpadlet I'm sorry man another one up: Solve the following equation for x: 3x^2 + 5x - 2 = 0
You're good. Utilize the quadratic formula. The quadratic formula states that for any equation in the form ax^2 + bx + c = 0, the solutions can be found utilizing the formula: x = (-b ± √(b^2 - 4ac)) / (2a) In this case, you have a = 3, b = 5, and c = -2. Plugging these values into the formula, you get: x = (-5 ± √(5^2 - 4*3*(-2))) / (2*3) After simplifying, you have: x = (-5 ± √(25 + 24)) / 6 x = (-5 ± √49) / 6 x = (-5 ± 7) / 6 Now you have two possibilities: 1. x = (-5 + 7) / 6 = 2 / 6 = 1/3 2. x = (-5 - 7) / 6 = -12 / 6 = -2 So the solutions to the equation are x = 1/3 and x = -2.
Bro idk how u do it
Good teacher.
So how is V?
She's pretty good. She typically hangs out on Johnson's Padlet.
makes sense
Yeah she tends to love Madi's Padlet as well.

Vocaloading:

@k1ngofpadlet wrote:
@vocaloading wrote:
@k1ngofpadlet wrote:
@vocaloading wrote:
@k1ngofpadlet wrote:
@vocaloading wrote:
@k1ngofpadlet wrote:
@vocaloading wrote:
@K1ngofpadlet I'm sorry man another one up: Solve the following equation for x: 3x^2 + 5x - 2 = 0
You're good. Utilize the quadratic formula. The quadratic formula states that for any equation in the form ax^2 + bx + c = 0, the solutions can be found utilizing the formula: x = (-b ± √(b^2 - 4ac)) / (2a) In this case, you have a = 3, b = 5, and c = -2. Plugging these values into the formula, you get: x = (-5 ± √(5^2 - 4*3*(-2))) / (2*3) After simplifying, you have: x = (-5 ± √(25 + 24)) / 6 x = (-5 ± √49) / 6 x = (-5 ± 7) / 6 Now you have two possibilities: 1. x = (-5 + 7) / 6 = 2 / 6 = 1/3 2. x = (-5 - 7) / 6 = -12 / 6 = -2 So the solutions to the equation are x = 1/3 and x = -2.
Bro idk how u do it
Good teacher.
So how is V?
She's pretty good. She typically hangs out on Johnson's Padlet.
makes sense
Yeah she tends to love Madi's Padlet as well.
Ikr? Madi has reallly stepped her game ALL THESE MF know her now

K1NGofPadlet:

@vocaloading wrote:
@k1ngofpadlet wrote:
@vocaloading wrote:
@k1ngofpadlet wrote:
@vocaloading wrote:
@k1ngofpadlet wrote:
@vocaloading wrote:
@k1ngofpadlet wrote:
@vocaloading wrote:
@K1ngofpadlet I'm sorry man another one up: Solve the following equation for x: 3x^2 + 5x - 2 = 0
You're good. Utilize the quadratic formula. The quadratic formula states that for any equation in the form ax^2 + bx + c = 0, the solutions can be found utilizing the formula: x = (-b ± √(b^2 - 4ac)) / (2a) In this case, you have a = 3, b = 5, and c = -2. Plugging these values into the formula, you get: x = (-5 ± √(5^2 - 4*3*(-2))) / (2*3) After simplifying, you have: x = (-5 ± √(25 + 24)) / 6 x = (-5 ± √49) / 6 x = (-5 ± 7) / 6 Now you have two possibilities: 1. x = (-5 + 7) / 6 = 2 / 6 = 1/3 2. x = (-5 - 7) / 6 = -12 / 6 = -2 So the solutions to the equation are x = 1/3 and x = -2.
Bro idk how u do it
Good teacher.
So how is V?
She's pretty good. She typically hangs out on Johnson's Padlet.
makes sense
Yeah she tends to love Madi's Padlet as well.
Ikr? Madi has reallly stepped her game ALL THESE MF know her now
Her Padlet is more popular than RNL

Vocaloading:

@k1ngofpadlet wrote:
@vocaloading wrote:
@k1ngofpadlet wrote:
@vocaloading wrote:
@k1ngofpadlet wrote:
@vocaloading wrote:
@k1ngofpadlet wrote:
@vocaloading wrote:
@k1ngofpadlet wrote:
@vocaloading wrote:
@K1ngofpadlet I'm sorry man another one up: Solve the following equation for x: 3x^2 + 5x - 2 = 0
You're good. Utilize the quadratic formula. The quadratic formula states that for any equation in the form ax^2 + bx + c = 0, the solutions can be found utilizing the formula: x = (-b ± √(b^2 - 4ac)) / (2a) In this case, you have a = 3, b = 5, and c = -2. Plugging these values into the formula, you get: x = (-5 ± √(5^2 - 4*3*(-2))) / (2*3) After simplifying, you have: x = (-5 ± √(25 + 24)) / 6 x = (-5 ± √49) / 6 x = (-5 ± 7) / 6 Now you have two possibilities: 1. x = (-5 + 7) / 6 = 2 / 6 = 1/3 2. x = (-5 - 7) / 6 = -12 / 6 = -2 So the solutions to the equation are x = 1/3 and x = -2.
Bro idk how u do it
Good teacher.
So how is V?
She's pretty good. She typically hangs out on Johnson's Padlet.
makes sense
Yeah she tends to love Madi's Padlet as well.
Ikr? Madi has reallly stepped her game ALL THESE MF know her now
Her Padlet is more popular than RNL
It rrally is tbh

Vocaloading:

@vocaloading wrote:
@k1ngofpadlet wrote:
@vocaloading wrote:
@k1ngofpadlet wrote:
@vocaloading wrote:
@k1ngofpadlet wrote:
@vocaloading wrote:
@k1ngofpadlet wrote:
@vocaloading wrote:
@k1ngofpadlet wrote:
@vocaloading wrote:
@K1ngofpadlet I'm sorry man another one up: Solve the following equation for x: 3x^2 + 5x - 2 = 0
You're good. Utilize the quadratic formula. The quadratic formula states that for any equation in the form ax^2 + bx + c = 0, the solutions can be found utilizing the formula: x = (-b ± √(b^2 - 4ac)) / (2a) In this case, you have a = 3, b = 5, and c = -2. Plugging these values into the formula, you get: x = (-5 ± √(5^2 - 4*3*(-2))) / (2*3) After simplifying, you have: x = (-5 ± √(25 + 24)) / 6 x = (-5 ± √49) / 6 x = (-5 ± 7) / 6 Now you have two possibilities: 1. x = (-5 + 7) / 6 = 2 / 6 = 1/3 2. x = (-5 - 7) / 6 = -12 / 6 = -2 So the solutions to the equation are x = 1/3 and x = -2.
Bro idk how u do it
Good teacher.
So how is V?
She's pretty good. She typically hangs out on Johnson's Padlet.
makes sense
Yeah she tends to love Madi's Padlet as well.
Ikr? Madi has reallly stepped her game ALL THESE MF know her now
Her Padlet is more popular than RNL
It rrally is tbh
Is that D3AD4YOU guy still bothering you?

BACKSTREETBOYS3:

Hey Voc want my snap?

Vocaloading:

Sure!

K1NGofPadlet:

@vocaloading wrote:
@vocaloading wrote:
@k1ngofpadlet wrote:
@vocaloading wrote:
@k1ngofpadlet wrote:
@vocaloading wrote:
@k1ngofpadlet wrote:
@vocaloading wrote:
@k1ngofpadlet wrote:
@vocaloading wrote:
@k1ngofpadlet wrote:
@vocaloading wrote:
@K1ngofpadlet I'm sorry man another one up: Solve the following equation for x: 3x^2 + 5x - 2 = 0
You're good. Utilize the quadratic formula. The quadratic formula states that for any equation in the form ax^2 + bx + c = 0, the solutions can be found utilizing the formula: x = (-b ± √(b^2 - 4ac)) / (2a) In this case, you have a = 3, b = 5, and c = -2. Plugging these values into the formula, you get: x = (-5 ± √(5^2 - 4*3*(-2))) / (2*3) After simplifying, you have: x = (-5 ± √(25 + 24)) / 6 x = (-5 ± √49) / 6 x = (-5 ± 7) / 6 Now you have two possibilities: 1. x = (-5 + 7) / 6 = 2 / 6 = 1/3 2. x = (-5 - 7) / 6 = -12 / 6 = -2 So the solutions to the equation are x = 1/3 and x = -2.
Bro idk how u do it
Good teacher.
So how is V?
She's pretty good. She typically hangs out on Johnson's Padlet.
makes sense
Yeah she tends to love Madi's Padlet as well.
Ikr? Madi has reallly stepped her game ALL THESE MF know her now
Her Padlet is more popular than RNL
It rrally is tbh
Is that D3AD4YOU guy still bothering you?
He's not worth talking about honestly.

Vocaloading:

@k1ngofpadlet wrote:
@vocaloading wrote:
@vocaloading wrote:
@k1ngofpadlet wrote:
@vocaloading wrote:
@k1ngofpadlet wrote:
@vocaloading wrote:
@k1ngofpadlet wrote:
@vocaloading wrote:
@k1ngofpadlet wrote:
@vocaloading wrote:
@k1ngofpadlet wrote:
@vocaloading wrote:
@K1ngofpadlet I'm sorry man another one up: Solve the following equation for x: 3x^2 + 5x - 2 = 0
You're good. Utilize the quadratic formula. The quadratic formula states that for any equation in the form ax^2 + bx + c = 0, the solutions can be found utilizing the formula: x = (-b ± √(b^2 - 4ac)) / (2a) In this case, you have a = 3, b = 5, and c = -2. Plugging these values into the formula, you get: x = (-5 ± √(5^2 - 4*3*(-2))) / (2*3) After simplifying, you have: x = (-5 ± √(25 + 24)) / 6 x = (-5 ± √49) / 6 x = (-5 ± 7) / 6 Now you have two possibilities: 1. x = (-5 + 7) / 6 = 2 / 6 = 1/3 2. x = (-5 - 7) / 6 = -12 / 6 = -2 So the solutions to the equation are x = 1/3 and x = -2.
Bro idk how u do it
Good teacher.
So how is V?
She's pretty good. She typically hangs out on Johnson's Padlet.
makes sense
Yeah she tends to love Madi's Padlet as well.
Ikr? Madi has reallly stepped her game ALL THESE MF know her now
Her Padlet is more popular than RNL
It rrally is tbh
Is that D3AD4YOU guy still bothering you?
He's not worth talking about honestly.
damn

K1NGofPadlet:

@vocaloading wrote:
@k1ngofpadlet wrote:
@vocaloading wrote:
@vocaloading wrote:
@k1ngofpadlet wrote:
@vocaloading wrote:
@k1ngofpadlet wrote:
@vocaloading wrote:
@k1ngofpadlet wrote:
@vocaloading wrote:
@k1ngofpadlet wrote:
@vocaloading wrote:
@k1ngofpadlet wrote:
@vocaloading wrote:
@K1ngofpadlet I'm sorry man another one up: Solve the following equation for x: 3x^2 + 5x - 2 = 0
You're good. Utilize the quadratic formula. The quadratic formula states that for any equation in the form ax^2 + bx + c = 0, the solutions can be found utilizing the formula: x = (-b ± √(b^2 - 4ac)) / (2a) In this case, you have a = 3, b = 5, and c = -2. Plugging these values into the formula, you get: x = (-5 ± √(5^2 - 4*3*(-2))) / (2*3) After simplifying, you have: x = (-5 ± √(25 + 24)) / 6 x = (-5 ± √49) / 6 x = (-5 ± 7) / 6 Now you have two possibilities: 1. x = (-5 + 7) / 6 = 2 / 6 = 1/3 2. x = (-5 - 7) / 6 = -12 / 6 = -2 So the solutions to the equation are x = 1/3 and x = -2.
Bro idk how u do it
Good teacher.
So how is V?
She's pretty good. She typically hangs out on Johnson's Padlet.
makes sense
Yeah she tends to love Madi's Padlet as well.
Ikr? Madi has reallly stepped her game ALL THESE MF know her now
Her Padlet is more popular than RNL
It rrally is tbh
Is that D3AD4YOU guy still bothering you?
He's not worth talking about honestly.
damn
Bro's not even worth taking a shit over.

Vocaloading:

@k1ngofpadlet wrote:
@vocaloading wrote:
@k1ngofpadlet wrote:
@vocaloading wrote:
@vocaloading wrote:
@k1ngofpadlet wrote:
@vocaloading wrote:
@k1ngofpadlet wrote:
@vocaloading wrote:
@k1ngofpadlet wrote:
@vocaloading wrote:
@k1ngofpadlet wrote:
@vocaloading wrote:
@k1ngofpadlet wrote:
@vocaloading wrote:
@K1ngofpadlet I'm sorry man another one up: Solve the following equation for x: 3x^2 + 5x - 2 = 0
You're good. Utilize the quadratic formula. The quadratic formula states that for any equation in the form ax^2 + bx + c = 0, the solutions can be found utilizing the formula: x = (-b ± √(b^2 - 4ac)) / (2a) In this case, you have a = 3, b = 5, and c = -2. Plugging these values into the formula, you get: x = (-5 ± √(5^2 - 4*3*(-2))) / (2*3) After simplifying, you have: x = (-5 ± √(25 + 24)) / 6 x = (-5 ± √49) / 6 x = (-5 ± 7) / 6 Now you have two possibilities: 1. x = (-5 + 7) / 6 = 2 / 6 = 1/3 2. x = (-5 - 7) / 6 = -12 / 6 = -2 So the solutions to the equation are x = 1/3 and x = -2.
Bro idk how u do it
Good teacher.
So how is V?
She's pretty good. She typically hangs out on Johnson's Padlet.
makes sense
Yeah she tends to love Madi's Padlet as well.
Ikr? Madi has reallly stepped her game ALL THESE MF know her now
Her Padlet is more popular than RNL
It rrally is tbh
Is that D3AD4YOU guy still bothering you?
He's not worth talking about honestly.
damn
Bro's not even worth taking a shit over.
Damn homie roasting mateo

BACKSTREETBOYS3:

Voc

Vocaloading:

@backstreetboys3 wrote:
Ummm anyway here you add me
Got it thx

Vocaloading:

@littleangel222

K1NGofPadlet:

@vocaloading wrote:
@k1ngofpadlet wrote:
@vocaloading wrote:
@k1ngofpadlet wrote:
@vocaloading wrote:
@vocaloading wrote:
@k1ngofpadlet wrote:
@vocaloading wrote:
@k1ngofpadlet wrote:
@vocaloading wrote:
@k1ngofpadlet wrote:
@vocaloading wrote:
@k1ngofpadlet wrote:
@vocaloading wrote:
@k1ngofpadlet wrote:
@vocaloading wrote:
@K1ngofpadlet I'm sorry man another one up: Solve the following equation for x: 3x^2 + 5x - 2 = 0
You're good. Utilize the quadratic formula. The quadratic formula states that for any equation in the form ax^2 + bx + c = 0, the solutions can be found utilizing the formula: x = (-b ± √(b^2 - 4ac)) / (2a) In this case, you have a = 3, b = 5, and c = -2. Plugging these values into the formula, you get: x = (-5 ± √(5^2 - 4*3*(-2))) / (2*3) After simplifying, you have: x = (-5 ± √(25 + 24)) / 6 x = (-5 ± √49) / 6 x = (-5 ± 7) / 6 Now you have two possibilities: 1. x = (-5 + 7) / 6 = 2 / 6 = 1/3 2. x = (-5 - 7) / 6 = -12 / 6 = -2 So the solutions to the equation are x = 1/3 and x = -2.
Bro idk how u do it
Good teacher.
So how is V?
She's pretty good. She typically hangs out on Johnson's Padlet.
makes sense
Yeah she tends to love Madi's Padlet as well.
Ikr? Madi has reallly stepped her game ALL THESE MF know her now
Her Padlet is more popular than RNL
It rrally is tbh
Is that D3AD4YOU guy still bothering you?
He's not worth talking about honestly.
damn
Bro's not even worth taking a Cheese over.
Damn homie roasting mateo
Just stating factual info.

LittleAngel222:

ya'll i cant breath this is too funny lmaooo

LittleAngel222:

like i love ya'll your my fav people